Professor Weil:

In response to your invitation to come and talk I wrote the enclosed letter. After I wrote it I realized there was hardly a statement in it of which I was certain. If you are willing to read it as pure speculation I would appreciate that; if not — I am sure you have a waste basket handy.

Yours truly, R. Langlands

Jan. 16, ’67

Dear André,

Langlands gave it to me to have a look and thus pass it on to you. Since I am busy right now, I am giving it to you after just a casual glance. I hope you will let me have it back a little later.

Yours Sincerely Harish